# BitMark for Infinity: Watermarking Bitwise Autoregressive Image Generative Models

### v1

@@ 4 Empirical Evaluation
Experimental Setup. We employ the Infinity-2B checkpoints and code  [10] to evaluate BitMark across different watermarking strengths and against a wide variety of attacks.  [^1] Infinity-2B was released six months ago, at the time of submission, allowing sufficient time for experimentation. Note that the most powerful Infinity-20B model (that we are aware of) is kept private (while the Infinity-8B was released less than 3 months ago before the submission, limiting our ability to test it). The Infinity-2B model generates $1024\times 1024$ pixel RGB-images and we choose the visual-tokenizer with a vocabulary size of $2^{32}$ per token. We follow [10] in the choice of hyper-parameters and use a classifier-free-guidance of 4, with a generation over 13 scales, following the scale schedule for an aspect ratio of 1:1. Hence, we utilize the best performing model, tokenizer and hyperparameters reported in [10]. We ablate the choice of the green, red list split for our experiments in Appendix B.
@@ 4.1 BitMark Introduces Detectable Watermarks while Maintaining Generation Performance
We demonstrate that our BitMark preserves the utility of the Infinity model with negligible impact on the image quality and the inference speed.
High Image Quality. Table 1 presents a quantitative evaluation of image generation quality for different watermarking strengths dependent on the logit bias parameter $\delta$. We generated 10,000 images based on prompts of the MS-COCO 2014 [16] validation dataset. We report the FID, KID, measuring the similarity of generated images and natural images in feature space, and the CLIP Score, which measures the alignment between prompt and image, for the varying watermark strengths $\delta$.
@@
The results show that, for $\delta\leq 3$, BitMark has a negligible impact on image quality (as measured by FID and KID) and minimally affects the prompt-following capabilities of the underlying model, as indicated by the CLIP Score. We also present a qualitative analysis of image quality in Section E.2, which shows images generated for different $\delta$. While there is an inherent trade-off in terms of watermark strength an image quality, we choose a small $\delta=2$ for our watermarks to obtain minimal impact on the generation while having a significantly strong watermark.
Fast Watermark Generation and Verification. For a watermark to be reasonably applicable in a real world setting, the watermark should have negligible impact on image generation speed [8, 34, 43]. We thus analyze the practicality of applying BitMark in a standard setting and its impact on generation speed of the Infinity model. Next to the generation performance, the speed of watermark detection has a strong impact on the possible applications, especially in a setting where we want to quickly detect data generated by Infinity. The quantitative analysis of BitMark in Section E.3 shows that generating a watermarked image requires at most one-tenth the time longer than the standard image generation. Verification, i.e., encoding an image into its bit representation and computing the $z$-score, takes less than 0.5 seconds per image (usually around 5X faster than the image generation).
@@ 4.2 BitMark is Robust Against a Wide Range of Attacks
Our results demonstrate that BitMark is extremely robust against a wide range of attacks, including conventional attacks, reconstruction attacks, watermarks in the sand attack, and even our own adaptive bit-flipping attack, which we carefully design with the aim of exploiting knowledge of our watermark method to remove BitMark.
@@
Robustness against Conventional Attacks. We perform a wide range of conventional attacks, namely: 1) Noise: Adds Gaussian noise with a std of 0.1 to the image, 2) Blur Noise attack + application of a Gaussian blur with a kernel size of 7, 3) Color: Color jitter with a random hue of 0.3, saturation scaling of 3.0 and contrast of 3.0, 4) Crop: Random crop to 70% 5) Rotation: Random rotation between different degrees in Table 2 and 0° and 180° in Table 3. 5) JPEG: 25% JPEG compression. Table 3 shows that our approach achieves a high TPR@1%FPR for most conventional attacks. Furthermore, BitMark achieves a high TPR@1%FPR against most attacks even for a low watermarking strength $\delta=1$. Although BitMark has slightly lower performance towards arbitrary rotation angles, the watermark is still highly detectable and it still obtains high TPR@1%FPR within a given rotation threshold. We report TPR@1%FPR and mean (std) of $z$ for different degrees of rotations for 1,000 images in Section E.5. Notably, it achieves as high as 99.9% TPR@1%FPR when the rotation range is between -20° and 20° Table 2. Given the strength of off-the-shelf tools for rotation estimation [19], it is highly achievable to rotate the image back into a tolerable range of rotations during the pre-processing stage in the model training process, thus ensuring that the watermark is robust against image rotations.
Robustness against Reconstruction Attacks. We also evaluate our proposed method against two state-of-the-art reconstruction attacks (shown as the last two columns in Table 3): 1) SD2.1-VAE: Encoding and decoding with the Stable Diffusion 2.1 autoencoder [24], 2) CtrlRegen+: Application of controllable regeneration from noise in the latent space following  [18]. For the reconstruction attacks, we find that the strength $\delta=2$ can withstand the CtrlRegen+ reconstruction attack, which removes watermarks such as TreeRing [34] or StableSignature [8].
Table 3: BitMark is robust against watermark removal attacks. We report the TPR@1%FPR (%) for the different attacks.
$\delta$ Conventional Attacks Reconstruction Attacks
None Noise Blur Color Crop Rotation JPEG SD2.1-VAE CtrlRegen+
1 100.0 97.8 99.2 99.3 68.5 14.5 100.0 100.0 61.9
2 100.0 99.6 99.9 99.8 98.8 20.1 100.0 100.0 91.6
3 100.0 99.7 99.9 99.9 99.8 23.8 100.0 100.0 97.0
4 100.0 99.9 99.9 100.0 99.9 29.2 100.0 100.0 99.0
5 100.0 100.0 100.0 99.9 100.0 34.3 100.0 100.0 99.8
@@
Watermarked Image, before the attack 100.0 100.0 100.0 57.6 87.8 59.6
Watermarked Image, after the attack 98.5 99.8 89.0 50.9 10.8 6.8
Robustness against Watermarks in the Sand Attack. We also evaluate the robustness of our watermarking scheme against the attack proposed in [38], namely, the watermarks in the sand attack. The attack leverages the CLIP Score as a quality oracle, the stable-diffusion-2-base [21] as a perturbation oracle, and iteratively modifies the watermarked image to remove the watermark. During each iteration, the attack produces potential candidates by inpainting randomly masked regions in the image. The attack successfully replaces most visual details in the original watermarked image while selecting the candidate with the best CLIP Score. As a result, many watermarks for diffusion models (e.g., [8]) can be removed through this attack. However, results in Table 4 show that our proposed method achieves a significant detection rate with TPR@1%FPR of 89.0%. Although a large percentage of the green-list patterns are removed, the average $z$-score in the attacked images is still as high as 10.8, which remains significant when compared with a mean $z$-score of 0.2 for non-watermarked images. Moreover, the attack causes a quality loss of the image when compared to the original image generated by Infinity. With 200 iterations and 100 candidates during each iteration, the method further requires a high computational cost, where more than 12 minutes are needed to attack a single image in the environment specified in Section C.2. We note that, the results do not conclude the failure of the theoretical framework in watermarks in the sand, but rather that there are currently no successful instantiations of the perturbation oracles and quality oracle for Infinity that can remove our proposed watermark. A more detailed analysis about this attack can be found in Appendix F.
@@
Robustness Against our Novel Bit-Flipper Attack. We further demonstrate that, even giving the attacker much stronger knowledge than our threat model formulated in Section 3, BitMark cannot be erased with severe image degradation. Concretely, we assume the attacker is given the knowledge of the auto-encoder of the generative model, the watermarking embedding and detection algorithms, and even our red and green lists. We propose a novel Bit-Flipper attack tailored for this broader attack surface, aiming at removing our watermarks. During this attack, the attacker first computes the total number of bit sequences $|s|$ and calculate the green fraction $C/|s|$. Then, the attacker randomly flips bits to convert green bit sequences to red ones. The goal of the bit flipping is to reduce the green fraction to the level of non-watermarked images, which is 50%, as shown in Appendix D. Specifically, the attacker computes the probability of flipping a bit via a flip factor $\phi$, such that $p=($$C/|s|$$)-0.5)*\phi$, whereas $\phi$ determines the bit sequence removal severity. A detailed algorithm of our attack is described in Algorithm 3. As shown in Appendix G, $\phi=2.2$ has the most impact on detectability. However, randomly flipping bits comes with very high degradation in image quality, as shown in Section E.2. Considering a symmetric detection threshold $|z|$, if we remove too many bit sequences in $G$, this further leads to the detectability of the watermark (see Appendix G).
@@ 4.3 Our BitMark Watermark Exhibits High Radioactivity
Training Setup. Given the extensive compute costs of training high-performant image-autoregressive models from scratch, we simulate full training by full fine-tuning the models. Therefore, we first generate 1,000 images with prompts of the MS-COCO2014 [16] validation dataset with our BitMark embedded under a watermarking strength of $\delta=2$. The watermarked outputs of $M_{1}$ are then used to train (i.e., in our approximative setup full fine-tune) the $M_{2}$ model. We analyze two different settings, with the same $M_{1}$ of Infinity-2B, namely: (1) the same modality setting, transfer to Infinity-2B and (2) the cross-modality setting, transfer to the Latent Diffusion Model (LDM) Stable Diffusion 2.1 [24]. To test the radioactivity, we compute the TPR@1%FPR of output images of $M_{2}$ and report them in Table 5.
@@
BitMark is Highly Radioactive. Table 5 highlights the strong radioactivity of our BitMark. When fine-tuning a model $M_{2}$ of the same architecture, the watermark is still detectable with a 99.7% TPR@1%FPR. Even in the cross-modality setting, BitMark is still detectable with a 98.9%TPR@1%FPR. Dubiński et al. [6] attribute the lack of radioactivity of watermarks in LDMs to the encoding from the pixel space into the latent space and the noising-denoising process while training $M_{2}$. However, our BitMark remains detectable despite the encoding and the noising-denoising process. The robustness of our watermark arises from slightly changing all parts of a given generated image, including low-level and high-level features, which are then also partly adapted by $M_{2}$. The observed results are in line with our robustness against CtrlRegen+ (see Table 3), an attack which simulates noising and denoising in the latent space. BitMark’s radioactivity is a very useful property that allows model owners to track provenance of their data and data generated based on it, even when this new data does not originate directly from their original model.
@@ 5 Conclusions
We present BitMark, the first bitwise watermarking framework for image autoregressive models, such as Infinity. We show that our watermark can be efficiently embedded and detected, imposes negligible overhead on image quality and generation speed, while exhibiting a strong robustness against a wide range of watermark removal attacks, including the strong watermarks in the sand and our own adaptive Bit-Flipper attack. Moreover, BitMark exhibits strong radioactivity, i.e., downstream models fine-tuned on our watermarked outputs consistently inherit and reproduce our bit patterns, which makes our watermark reliably detectable. This is important for model owners to identify images derived from their own generated images, and exclude them from training of future model versions. Overall, BitMark not only sets a new standard for watermarking in autoregressive image generation, but also provides a critical defense against model collapse, contributing to further advancements in image generative models.
@@
This work was supported by the German Research Foundation (DFG) within the framework of the Weave Programme under the project titled "Protecting Creativity: On the Way to Safe Generative Models" with number 545047250. Responsibility for the content of this publication lies with the authors.
Appendix A Additional Algorithms
We report the functions used in Algorithm 1, Algorithm 2 and Algorithm 3 in detail here:

### v2

@@ 4 Empirical Evaluation
Experimental Setup. In the following, we employ the Infinity-2B checkpoint [11] as the default model to evaluate BitMark across different watermarking strengths and against a wide variety of attacks. Additionally we show the generalizability of BitMark to other models, namely the Infinity-8B checkpoint as well as the Instella-T2I [37] IAR model. We further show generalizability of our method to bitwise diffusion models in Appendix˜G. The Infinity models natively generate $1024\times 1024$ pixel RGB-images, we choose the visual-tokenizer with a vocabulary size of $2^{32}$ and $2^{56}$ per token for the Infinity-2B and the Infinity-8B checkpoint, respectively. We follow Han et al. [11] in the choice of hyper-parameters and use a classifier-free-guidance of 4, with a generation over 13 scales, following the scale schedule for an aspect ratio of 1:1. Hence, we utilize the best performing model, tokenizer and hyperparameters reported by Han et al. [11]. We ablate the choice of the green and red list split for our experiments in Appendix˜D. For Instella IAR we follow Wang et al. [37], leveraging a vocabulary of size $2^{32}$ and a cfg of 7.5 at a $1024\times 1024$ resolution. We evaluate our experiments based on images and prompts of the MS-COCO 2014 [18] validation dataset. Unless stated otherwise, we use a subset of 10,000 images.
Baselines. To the best our knowledge no prior work developed an in generation-time watermarking scheme for autoregressive image models. Therefore we compare the robustness of BitMark against three postprocessing watermarks in Table˜2. Namely we analyze RivaGAN [43] (with message length = 32), StegaStamp [34] (with message length = 100) and TrustMark [5] (with message length = 100). We applied the postprocessing watermarks on 1,000 images generated by Infinity-2B.
@@ 4.1 BitMark Introduces Detectable Watermarks while Maintaining Generation Performance
We demonstrate that our BitMark preserves the utility of the model with negligible impact on the image quality and the inference speed.
High Image Quality. Table˜1 presents a quantitative evaluation of image generation quality for different watermarking strengths dependent on the logit bias parameter $\delta$. We report the FID, KID, measuring the similarity of generated images and natural images in feature space, and the CLIP Score, which measures the alignment between prompt and image, for the varying watermark strengths $\delta$.
@@
The results show that, for $\delta\leq 3$, BitMark has a negligible impact on image quality (as measured by FID and KID) and minimally affects the prompt-following capabilities of the underlying model, as indicated by the CLIP Score. We also present a qualitative analysis of image quality in Section˜F.3, which shows images generated for different $\delta$. While there is an inherent trade-off in terms of watermark strength an image quality, we choose a small $\delta=2$ for the Infinity-2B checkpoint to obtain minimal impact on the generation while having a significantly strong watermark. We further show that $\delta=1.5$ does not negatively impact image quality for Infintiy-8B and Instella in Table˜15.
Fast Watermark Generation and Verification. For a watermark to be reasonably applicable in a real world setting, the watermark should also have negligible impact on image generation speed [9, 38, 48]. We thus analyze the practicality of applying BitMark in a standard setting and its impact on generation speed of the Infinity model. Next to the generation performance, the speed of watermark detection has a strong impact on the possible applications, especially in a setting where we want to quickly detect generated data. The quantitative analysis of BitMark in Section˜F.4 shows that generating a watermarked image requires at most one-tenth the time longer than the standard image generation. Verification, i.e., encoding an image into its bit representation and computing the $z$-score, takes less than 0.5 seconds per image, usually around $5\times$ faster than the image generation.
@@ 4.2 BitMark is Robust Against a Wide Range of Attacks
Our results demonstrate that BitMark is robust against a wide range of attacks, including conventional attacks, reconstruction attacks, watermarks in the sand attack, and our own adaptive bit-flipping attack, with the aim of exploiting knowledge of our watermark method to remove BitMark.
Robustness against Conventional Attacks. We perform a wide range of conventional attacks, namely: 1) Noise: Adds Gaussian noise with a std of 0.1 to the image, 2) Blur Noise attack + application of a Gaussian blur with a kernel size of 7, 3) Color: Color jitter, a random hue of 0.3, saturation scaling of 3.0 and contrast of 3.0, 4) Crop: Random crop to 70%, 5) Rotation: Random rotation between 0° and 180°, 6) JPEG: 25% JPEG compression, 7) Horizontal Flipping, 8) Vertical Flipping.
Table 2: BitMark is robust against watermark removal attacks. We report the TPR@1%FPR (%) for the different conventional and reconstruction attacks.
Watermark Conventional Attacks Reconstruction Attacks
None Noise Blur Color Rotate Crop JPEG Vertical Horizontal SD2.1-VAE CtrlRegen+
RivaGAN [43] 99.7 98.3 99.7 99.4 96.7 99.4 99.7 0.0 0.0 98.5 1.6
StegaStamp [34] 100.0 100.0 100.0 98.7 32.1 1.0 100.0 1.0 33.8 100.0 44.2
TrustMark [5] 99.9 99.5 99.9 2.2 2.7 1.6 99.9 0.7 99.8 99.7 1.1
Infinity-2B ($\delta=2$) 100.0 99.6 99.9 99.8 20.1 98.8 100.0 78.8 100.0 100.0 91.6
Infinity-8B ($\delta=1.5$) 100.0 99.7 100.0 75.5 57.6 99.4 99.9 93.8 99.7 100.0 30.4
Instella IAR ($\delta=1.5$) 100.0 96.0 100.0 75.3 2.7 7.5 100.0 9.3 12.1 100.0 93.6
@@
Table˜2 shows that our approach achieves a high TPR@1%FPR for most conventional attacks compared to the baselines. Each baseline is weak against at least one conventional attacks. RivaGAN is not robustness against flipping (0%TPR@1%FPR), StegaStamp is not robustness against cropping (1%TPR@1%FPR) and TrustMark is vulnerable to color jitter (2.2%TPR@1%FPR), rotation (2.7%TPR@1%FPR) and cropping (1.6%TPR@1%FPR). Although BitMark for Infinity-2B has slightly lower performance towards arbitrary rotation angles, the watermark is still highly detectable and it obtains high TPR@1%FPR within a given rotation threshold. We report TPR@1%FPR and mean (std) of $z$ for different degrees of rotations for 1,000 images in Section˜F.2. Notably, it achieves as high as 99.9% TPR@1%FPR when the rotation range is between -20° and 20° as shown in  Table˜3. Given the strength of off-the-shelf tools for rotation estimation [21], it is highly achievable to rotate the image back into a tolerable range of rotations during the pre-processing stage of the detection, resulting in a TPR@1%FPR of 98.3% for randomly rotated images as shown in Table˜13. Table˜12 provides an ablation of the robustness of BitMark given different values of $\delta$, showing that BitMark for Infinity-2B is highly robust against most attacks even with $\delta=1$. Infinity-8B is more robust against rotation compared to the smaller 2B checkpoint. The effectiveness of BitMark on Instella IAR is by design very limited, given its total number of 128 tokens, corresponding to 4,096 bits, per image.
Robustness against Reconstruction Attacks. We also evaluate our proposed method against two state-of-the-art reconstruction attacks (shown as the last two columns in Table˜2): 1) SD2.1-VAE: Encoding and decoding with the Stable Diffusion 2.1 autoencoder [26], 2) CtrlRegen+: Application of controllable regeneration from noise in the latent space following  [20]. Regarding reconstruction attacks, our results indicate that BitMark for Infinity-2B is robust against the CtrlRegen+ reconstruction attack, achieving 91.6%TPR@1%FPR. Also Instella IAR achieves strong robustness of 93.6%TPR@1%FPR. In contrast, RivaGAN and TrustMark demonstrate a lack of robustness with 1.6%TPR@1%FPR and 1.1%TPR@1%FPR, respectively. StegaStamp exhibits partial robustness of 44.2%TPR@1%FPR. The Infinity-8B shows only partial robustness, which is further discussed in Appendix˜A.
@@
Watermarked Image (before attack) 100.0 100.0 100.0 57.6 87.8 59.6
Watermarked Image (after attack) 98.5 99.8 89.0 50.9 10.8 6.8
Robustness against Watermarks in the Sand Attack. We also evaluate the robustness of our watermarking scheme against the attack proposed in [42], namely, the watermarks in the sand attack. The attack leverages the CLIP Score as a quality oracle, the stable-diffusion-2-base [23] as a perturbation oracle, and iteratively modifies the watermarked image to remove the watermark. During each iteration, the attack produces potential candidates by inpainting randomly masked regions in the image. The attack successfully replaces most visual details in the original watermarked image while selecting the candidate with the best CLIP Score. As a result, many watermarks for diffusion models (e.g., [9]) can be removed through this attack. However, results in Table˜4 show that our proposed method achieves a significant detection rate with TPR@1%FPR of 89.0%. Although a large percentage of the green-list patterns are removed, the average $z$-score in the attacked images is still as high as 10.8, which remains significant when compared with a mean $z$-score of 0.2 for non-watermarked images. Moreover, the attack causes a quality loss of the image when compared to the original image generated by Infinity. With 200 iterations and 100 candidates during each iteration, the method further requires a high computational cost, where more than 12 minutes are needed to attack a single image in the environment specified in Section˜E.2. We note that the results do not conclude the failure of the theoretical framework in watermarks in the sand, but rather that there are currently no successful instantiations of the perturbation oracles and quality oracle for Infinity that can remove our proposed watermark. A more detailed analysis can be found in Section˜H.2.
@@
[…]
@@ 4.3 Our BitMark Watermark Exhibits High Radioactivity
Training Setup. Given the extensive compute costs of training high-performant image-autoregressive models from scratch, we simulate full training by full fine-tuning the models. Therefore, we first generate 1,000 images with prompts of the MS-COCO2014 [18] validation dataset with our BitMark embedded under a watermarking strength of $\delta=2$. The watermarked outputs of $M_{1}$ are then used to train (i.e., in our approximative setup full fine-tune) the $M_{2}$ model. We analyze two different settings, with the same $M_{1}$ of Infinity-2B, namely: (1) the same modality setting, transfer to Infinity-2B and (2) the cross-modality setting, transfer to the Latent Diffusion Model (LDM) Stable Diffusion 2.1 [26]. To test the radioactivity, we compute the TPR@1%FPR of output images of $M_{2}$ and report them in Table˜5.
BitMark is Highly Radioactive. Table˜5 highlights the strong radioactivity of our BitMark. When fine-tuning a model $M_{2}$ of the same architecture, the watermark is still detectable with a 99.7%TPR@1%FPR. Even in the cross-modality setting, BitMark is still detectable with a 98.9%TPR@1%FPR. Dubiński et al. [7] attribute the lack of radioactivity of watermarks in LDMs to the encoding from the pixel space into the latent space and the noising-denoising process while training $M_{2}$.
@@
However, our BitMark remains detectable despite the encoding and the noising-denoising process. The robustness of our watermark arises from slightly changing all parts of a given generated image, including low-level and high-level features, which are then also partly adapted by $M_{2}$. The observed results are in line with our robustness against CtrlRegen+ (see Table˜2), an attack which simulates noising and denoising in the latent space. BitMark’s radioactivity is a very useful property that allows model owners to track provenance of their data and data generated based on it, even when this new data does not originate directly from their original model.
We additionally ablate the ratio of watermarked to clean data needed for our BitMark to become radioactive to the Inifinity-2B model in Table˜6. Even when model $M_{2}$ is trained on only 10% of our watermarked data, it generates images with a clear bias towards our watermark distribution, resulting in a high TPR of 22.7% at 1% FPR. We note that even a slight bias is significant at the distribution level. To validate this claim, we apply a one-tailed Mann-Whitney U test to determine if the z-score distribution of $M_{2}$-generated data is significantly larger than on $M_{1}$ clean generated data. The test is significant with a p-value of 5.6e-62 when $M_{2}$ is trained on only 5% watermarked data, showing that our BitMark provides a traceable signal across training cycles in settings where only small fractions of the training data are watermarked.
Table 6: Detection performance and statistical significance across fractions of watermarked data. TPR@1%FPR (top row) and Mann–Whitney U-test p-values (bottom row) for different fractions of watermarked data used for fine-tuning.
[…]
TPR@1%FPR 1.0% 0.3% 1.0% 2.9% 22.7% 98.3% 100.0% 100.0%
Mann–Whitney U Test (p-value) 0.999 0.999 0.999 5.6e-62 6.8e-209 $\approx$ 0 $\approx$ 0 $\approx$ 0
@@ 5 Conclusions
We present BitMark, the first watermarking framework for bitwise image autoregressive models. We show that our watermark can be efficiently embedded and detected, imposes negligible overhead on image quality and generation speed, while exhibiting a strong robustness against a wide range of watermark removal attacks, including the strong watermarks in the sand and our own adaptive Bit-Flipper attack. Moreover, BitMark exhibits strong radioactivity, i.e., downstream models fine-tuned on our watermarked outputs consistently inherit and reproduce our bit patterns, which makes our watermark reliably detectable requiring only 5% watermarked data to achieve a significant distributional shift. This is important for model owners to identify images derived from their own generated images, and exclude them from training of future model versions. Overall, BitMark not only sets a new standard for watermarking in autoregressive image generation, but also provides a critical defense against model collapse, contributing to further advancements in image generative models.
@@
This work was supported by the German Research Foundation (DFG) within the framework of the Weave Programme under the project titled ”Protecting Creativity: On the Way to Safe Generative Models” with number 545047250. Responsibility for the content of this publication lies with the authors. We would like to also acknowledge our sponsors, who support our research with financial and in-kind contributions, especially the OpenAI Cybersecurity Grant.
Appendix A Limitations and Broader Impact
Limitations. While BitMark for Infinity-2B is robust against moderate rotations (99.9% TPR@1%FPR within rotation angles -20° to 20°), the detection accuracy degrades with larger rotation angles, dropping to 56.8% TPR@1%FPR for ±50° rotations. Although open-source methods have a strong capability for rotation correction, this can be a limitation for real-world scenarios. This rotation sensitivity, however, is an inherent issue in Infinity’s data augmentation pipeline, which makes the autoencoder less robust to large rotations. Moreover, the soft biasing parameter $\delta$ requires tuning, as values exceeding $\delta=3$ lead to noticeable quality degradation (FID increases from 29.61 to 127.44 when $\delta$ increases from 3 to 5), constraining the maximum achievable watermark strength. Although we evaluate the detection BitMark with a commonly adopted metric, i.e., true positive rate at very low false positive rates (1%), it could potentially become problematic in extreme-scale applications involving trillions of samples, where even small false positive rates might impact training data quality. It is observed that the Infinity-8B checkpoint shows increased sensitivity to larger values of $\delta$. We hypothesize that the lower $\delta$ chosen may account for its comparatively lower robustness relative to the Infinity-2B checkpoint, as illustrated in Table˜2.
Appendix B Additional Related Work
Appendix C Application of BitMark on Single-Scale models
We report the application of BitMark for single scale models in Algorithm˜3 and the detection in Algorithm˜4.
We report the functions used in Algorithm˜1, Algorithm˜2, Algorithm˜3 and Algorithm˜4 here: